# CAPE: Covariate-Adjusted Pre-Training for Epidemic Time Series Forecasting

### v3

@@ A.8 Full Results on Pre-train datasets
A.9 Full results for few-shot forecasting
We present the complete few-shot performance across different horizons in Table 11. While CAPE does not achieve state-of-the-art average performance on the ILI USA dataset with limited training data, it excels in short-term forecasting when the horizon is smaller.
@@ A.9 Full results for few-shot forecasting
A.10 Impact of the Environment Distribution and Pre-train Epochs
Figure 9: Downstream performance vs. environment distribution and pre-train epochs.
A.11 Impact of pre-train ratio on the downstream datasets
We provide additional evaluations for CAPE on downstream datasets to analyze the impact of the pre-training ratio. As shown in Figure LABEL:fig:pretrain_scale, increasing the pre-training ratio eventually improves downstream performance across all datasets.
@@ A.11 Impact of pre-train ratio on the downstream datasets
A.12 Mitigating Distribution Shift
Figure 11: CMD scores w/ and w/o environment estimate between train/test sets. We measure the CMD scores based on the last embedding output by our model.
A.13 Disentangling Disease-Specific Dynamics
After pre-training, both CAPE and PatchTST were frozen and applied to downstream datasets directly to generate latent space embeddings for each sample. We then employed the Davies-Bouldin Index (DBI) to assess the separability of these embeddings by disease. CAPE consistently achieved lower DBI scores (we show an example in Figure 16), demonstrating a superior ability to distinguish between different diseases in the latent space compared to PatchTST. This enhanced separability highlights the effectiveness of CAPE’s environment estimation and backdoor adjustment. These mechanisms are crucial for mitigating the confounding influence of noisy, spurious, and environment-dependent factors ($X_{sp}$), thereby enabling the model to better isolate and represent the unique, intrinsic, and causal disease dynamics ($X_{ca}$).
Figure 12: DBI between the embeddings of each pair of downstream datasets from the pre-trained model. (See Figure  16).
A.14 Visualization of the Estimated Environments
According to $\hat{\mathbf{e}}^{(l)}=\sum_{k=1}^{K}\mathbf{e}_{k}\pi_{k}^{(l)}$, an aggregated environment is the weighted sum of the learned latent environment representations. Therefore, the estimation shares the same latent space as the fixed representations and we are able to visualize them using t-SNE. As shown in Figure 13, we visualize the aggregated environments (Estimated) as well as the learned latent environment (Anchor) from a CAPE model with 8 environments.
@@ A.14 Visualization of the Estimated Environments
Figure 13: Visualization of the estimated environment representations using t-SNE.
A.15 Visualization of Distribution Shift for Downstream Datasets
@@
where:$\mu_{k}(X)$ denotes the $k$-th central moment of $X$, defined as: $\mu_{k}(X)=\mathbb{E}[(X-\mathbb{E}[X])^{k}],$ and similarly for $\mu_{k}(X_{\text{test}})$. $\|\cdot\|_{2}$ is the Euclidean norm. $K$ is the maximum order of moments considered. The CMD score aggregates the differences in the mean (first moment) and higher-order moments (e.g., variance, skewness), providing a robust measure of the distribution shift. In our experiments, we set $K=3$ to capture up to the third-order central moments. This score quantitatively complements the visual observations in Figure 14, offering a more comprehensive understanding of the distributional differences between training and test sets.
Impact of Distribution Shifts. Distribution shifts between training and test datasets pose significant challenges to the generalizability and robustness of predictive models. When the underlying data distributions differ, models trained on the training set may struggle to maintain their performance on the test set, leading to reduced accuracy and reliability. These discrepancies can arise from various factors, such as temporal changes in infection patterns or geographical variations. In this paper, we assume that the inherent infection pattern of a particular disease remains constant, and the distribution shifts for the disease are primarily caused by the rapidly changing environment, which results in diverse infection patterns. In the context of epidemic modeling, such shifts are especially critical, as they can undermine the model’s ability to accurately predict future infection trends, which is essential for effective public health interventions.
@@
A.16 Latent space visualization of Measle and Covid datasets from pre-trained models.
In order to demonstrate that CAPE effectively disentangles the underlying dynamics of diseases from the influence of the environment, we visualize the output embeddings for the Measles and COVID datasets by projecting them into a two-dimensional space using t-SNE. Specifically, we utilize the pre-trained model without fine-tuning on these two downstream datasets and visualize $\mathbf{x}^{(L)}$, the final-layer embeddings, as individual data points in the figure. As shown in Figure 16, CAPE (left) visually separates the two datasets more effectively than the pre-trained PatchTST model (right). To quantitatively evaluate the separability of the embeddings, we compute the Davies–Bouldin Index (DBI), which is defined as:
@@
where $K$ is the number of clusters (in this case, two: Measles and COVID), $\mu_{i}$ is the centroid of cluster $i$, $\sigma_{i}$ is the average intra-cluster distance for cluster $i$, defined as: $\sigma_{i}=\frac{1}{|C_{i}|}\sum_{\mathbf{x}\in C_{i}}\|\mathbf{x}-\mu_{i}\|$, where $C_{i}$ is the set of points in cluster $i$, $\|\mu_{i}-\mu_{j}\|$ is the Euclidean distance between the centroids of clusters $i$ and $j$. The DBI measures the ratio of intra-cluster dispersion to inter-cluster separation. Lower DBI values indicate better separability. As shown in Figure 16, CAPE achieves a significantly lower DBI compared to PatchTST, confirming its superior ability to disentangle the underlying disease dynamics from environmental factors. A more complete result is shown in Figure 12.
A.17 Hyper-parameter Sensitivity Analysis
While we treat $\hat{\mathbf{E}}_{j,c}^{(l)}=\hat{\mathbf{e}}^{(l)}_{c}\odot\mathbf{h}^{(l)}_{c}$ in the contrastive loss and aligned the combined representation of the input and environment, we further applied the contrastive loss on the aggregated environment representations at each layer, denoted as $\tilde{{\mathbf{E}}}_{j,c}^{(l)}=\hat{\mathbf{e}}^{(l)}_{c}$ and assigned a hyper-parameter $\beta$ to control for its weight:
$\mathcal{L}_{\text{final}}$ $=\sum_{\mathbf{x}\in\mathbf{X}}\mathcal{L}_{\text{recon}}(\mathbf{x},\hat{\mathbf{x}})+\alpha\,\mathcal{L}_{\text{CL}}(\hat{\mathbf{E}}^{(L)},\hat{\mathbf{E}^{\prime}}^{(L)})+\frac{\beta}{L}\sum_{l}\mathcal{L}_{\text{CL}}(\tilde{{\mathbf{E}}}^{(l)},\tilde{{\mathbf{E}}^{\prime}}^{(l)}),$ (23)
$\hskip 85.35826pt\mathbf{X}\sim\mathcal{D}^{\prime}_{s}\in\mathcal{D}_{\text{pre}}$
where $L$ denotes the number of layers. Using Equation (23), we further fine-tune our pre-trained model by varying $\alpha$ and $\beta$ across the values [1e-3, 1e-2, 1e-1, 1, 10]. The results of this sensitivity analysis are presented in Figure 15.
(a) ILI USA
(b) Measles
(c) ILI Japan
Figure 15: Hyperparameters sensitivity of $\alpha$ and $\beta$.
Figure 16: Output latent space of two pre-trained models without fine-tuning from Measle and Covid datasets. Left: CAPE; Right: Pre-trained PatchTST.

### v4

@@ A.9 Full Results on Pre-train datasets
A.10 Full results for few-shot forecasting
We present the complete few-shot performance across different horizons in Table 10. While CAPE does not achieve state-of-the-art average performance on the ILI USA dataset with limited training data, it excels in short-term forecasting when the horizon is smaller. Since the authors of PEM (Kamarthi & Prakash, 2023) did not release full code, we implemented the method to the best of our ability based on the paper description.
@@ A.10 Full results for few-shot forecasting
A.11 Impact of the compartment Distribution and Pre-train Epochs
Figure 9: Downstream performance vs. compartment distribution and pre-train epochs.
A.12 Impact of pre-training ratio on the downstream datasets
We provide additional evaluations for CAPE on downstream datasets to analyze the impact of the pre-training ratio. As shown in Figure 10, increasing the pre-training ratio eventually improves downstream performance across all datasets.
@@ A.12 Impact of pre-training ratio on the downstream datasets
A.13 Mitigating Distribution Shift
Figure 11: CMD scores w/ and w/o compartment estimate between train/test sets. We measure the CMD scores based on the last embedding output by our model.
A.14 Disentangling Disease-Specific Dynamics
After pre-training, both CAPE and PatchTST were frozen and applied to downstream datasets directly to generate latent space embeddings for each sample. We then employed the Davies-Bouldin Index (DBI) to assess the separability of these embeddings by disease. CAPE consistently achieved lower DBI scores (we show an example in Figure 15), demonstrating a superior ability to distinguish between different diseases in the latent space compared to PatchTST. This enhanced separability highlights the effectiveness of CAPE’s compartment estimation and backdoor adjustment. These mechanisms are crucial for mitigating the confounding influence of noisy, spurious, and compartment-dependent factors ($X_{sp}$), thereby enabling the model to better isolate and represent the unique, intrinsic, and causal disease dynamics ($X_{ca}$).
Figure 12: DBI between the embeddings of each pair of downstream datasets from the pre-trained model. (See Figure  15).
A.15 Visualization of the Estimated compartments
According to $\hat{\mathbf{e}}^{(l)}=\sum_{k=1}^{K}\mathbf{e}_{k}\pi_{k}^{(l)}$, an aggregated compartment is the weighted sum of the learned latent compartment representations. Therefore, the estimation shares the same latent space as the fixed representations and we are able to visualize them using t-SNE. As shown in Figure 13, we visualize the aggregated compartments (Estimated) as well as the learned latent compartment (Anchor) from a CAPE model with 8 compartments.
@@ A.15 Visualization of the Estimated compartments
Figure 13: Visualization of the estimated compartment representations using t-SNE.
A.16 Visualization of Distribution Shift for Downstream Datasets
@@
[…]
Impact of Distribution Shifts. Distribution shifts between training and test datasets pose significant challenges to the generalizability and robustness of predictive models. When the underlying data distributions differ, models trained on the training set may struggle to maintain their performance on the test set, leading to reduced accuracy and reliability. These discrepancies can arise from various factors, such as temporal changes in infection patterns or geographical variations. In this paper, we assume that the inherent infection pattern of a particular disease remains constant, and the distribution shifts for the disease are primarily caused by the rapidly changing compartment, which results in diverse infection patterns. In the context of epidemic modeling, such shifts are especially critical, as they can undermine the model’s ability to accurately predict future infection trends, which is essential for effective public health interventions.
@@
A.17 Latent space visualization of Measle and Covid datasets from pre-trained models.
In order to demonstrate that CAPE effectively disentangles the underlying dynamics of diseases from the influence of the compartment, we visualize the output embeddings for the Measles and COVID datasets by projecting them into a two-dimensional space using t-SNE. Specifically, we utilize the pre-trained model without fine-tuning on these two downstream datasets and visualize $\mathbf{x}^{(L)}$, the final-layer embeddings, as individual data points in the figure. As shown in Figure 15, CAPE (left) visually separates the two datasets more effectively than the pre-trained PatchTST model (right). To quantitatively evaluate the separability of the embeddings, we compute the Davies–Bouldin Index (DBI), which is defined as:
@@
where $K$ is the number of clusters (in this case, two: Measles and COVID), $\mu_{i}$ is the centroid of cluster $i$, $\sigma_{i}$ is the average intra-cluster distance for cluster $i$, defined as: $\sigma_{i}=\frac{1}{|C_{i}|}\sum_{\mathbf{x}\in C_{i}}\|\mathbf{x}-\mu_{i}\|$, where $C_{i}$ is the set of points in cluster $i$, $\|\mu_{i}-\mu_{j}\|$ is the Euclidean distance between the centroids of clusters $i$ and $j$. The DBI measures the ratio of intra-cluster dispersion to inter-cluster separation. Lower DBI values indicate better separability. As shown in Figure 15, CAPE achieves a significantly lower DBI compared to PatchTST, confirming its superior ability to disentangle the underlying disease dynamics from compartmental factors. A more complete result is shown in Figure 12.
Figure 15: Output latent space of two pre-trained models without fine-tuning from Measle and Covid datasets. Upper: CAPE; Lower: Pre-trained PatchTST.